OpenStudy (anonymous):

FAN & MEDAL!! 2x – 5y = –6 2x – 7y = –14 Select one: a. (4, 7) b. (7, 4) c. (8, –4) d. (–7, 2)

OpenStudy (anonymous):

Just plug in the numbers given as an answer and see if it is correct. for example if we take the firs equation 2X-5Y=-6 and substitute (x,y) with the first option a.(4,7) then we would have 2(4)-5(7)=-6 which is = to 8-35=-6 which is not true because 8-35 is -27 not -6 so can you check the other answers?